plzz solve this
At what rate can money be invested if $2,330 is equivalent to $2,547.86 paid 188 days later? (Do not round the intermediate calculations. Round your final answer to the nearest 0.01%.)

Answer by rothauserc(4718)

we use the following formula
A = Pe^rt, where A is the amount after t years for principal P at rate r
:
note that t = 188/365 for this problem
:
2547.86 = 2330 * e^(r*(188/365))
:
e^(r*(188/365)) = 2547.86 / 2330 = 1.093502146
:
use definition of logarithms
:
(r*(188/365)) = ln(1.093502146)
r = (365 * ln(1.093502146)) / 188
r = 0.173541043
**************************************************************
therefore the rate is 17.35%